# Precipitation Downscaling with Spatiotemporal Video Diffusion

### v2

@@ 2.1.3 Loss Function
To train our model, we use the angular parametrization suggested by [49]. More concretely, this results in the diffusion loss of the form
@@
where $\epsilon\sim\mathcal{N}(0,I)$, $n$ is sampled uniformly from $\{1,2,\dots,N\}$, and the sequences $\mathbf{x}^{0:T}$, $\mathbf{y}^{0:T}$ are sampled from the training distribution. Note that $\mathbf{c}=(\mathbf{x}^{0:T},\overline{\mathbf{y}}^{0:T})$ where $\bar{\mathbf{y}}^{0:T}=\mu_{\phi}\left(\mathbf{x}^{0:T}\right)$. Here, the scalars $\alpha_{n}^{2}=\prod_{i=1}^{n}(1-\beta_{i})$ and $\sigma_{n}^{2}=1-\alpha_{n}^{2}$ are used to define $\mathbf{v}\equiv\alpha_{n}\epsilon-\sigma_{n}\mathbf{r}_{0}^{0:T}$.
Training and inference are concurrent across multiple frames by the inherent mechanism of spatio-temporal attention. Algorithm 1 and 2 demonstrate the training and sampling strategy under the angular parametrization. We use DDIM sampling [52] to generate frame residuals with fewer diffusion steps.
@@ 2.1.4 Network Architecture
Both the deterministic downscaler and the conditional diffusion model employ a UNet backbone (refer to Appendix A.1 for implementation details). The downscaler processes the multi-channel input frames ($\mathbf{x}^{0:T}$), yielding an initial estimate ($\bar{\mathbf{y}}^{0:T}$) focusing solely on the precipitation channel. The diffusion UNet, featuring similar architecture but with key adaptations, conditions on diffusion step $n$ and concatenates feature maps from the downscaler with its own before every contraction in the UNet path. The channel concatenated input to the diffusion UNet (${\mathbf{x}}^{0:T}$, $\bar{\mathbf{y}}^{0:T}$, and $\mathbf{r}^{0:T}_{n}$), along with the conditioning variables, i.e., diffusion step $n$ and the feature maps from downscaler, yields the output $\mathbf{v}$.
Traditionally, most image diffusion models utilise a UNet architecture with convolutional residual blocks and transformer blocks [41]. Each transformer block contains a spatial self-attention and cross-attention. The spatial self-attention focuses on pixel correlations within feature maps, while cross-attention considers relationships between pixels and conditional inputs.
Temporal coherence in video data can be achieved by employing full attention across the entire three-dimensional data cube, encompassing spatial and temporal dimensions. However, the computational demands associated with this approach render it impractical for processing large sequences or high-resolution inputs. To optimize efficiency, we make several modifications. These include decoupling attention between spatial and temporal dimensions, utilizing a linear variant of self-attention [21] for non-bottleneck layers (where the effective number of “tokens” for attention is relatively large), focusing spatial attention on localized patches (instead of the entire feature map, which could be wasteful), and calculating per-channel temporal attention in large spatial dimensions (namely, the ultimate and penultimate expansion and contraction layers of UNet). These modifications dramatically reduce the time complexity and memory footprint of these transformer blocks. At each temporal attention block, the time dimension in the feature map tensor is ungrouped to facilitate inter-frame attention.
@@ 3 Experiments
Figure 3: Distributions of the fine-grid three-hourly average precipitation, including all gridpoints around the globe. The Swin-IR baseline overestimates large precipitation events, whereas all other baselines underestimate key extreme and rare precipitation events. Our model aligns best with the fine-grid ground truth than any the other model. This is also evident with the the EMD and PE metrics discussed in Tab. 1 and Sec. 3.4.
@@
We conduct a comprehensive evaluation of our proposed method, Spatiotemporal Video Diffusion (STVD), against five contemporary state-of-the-art baselines. The first two baselines are image super-resolution models based on the Swin Vision Transformer (Swin-IR)[30] and its residual diffusion variant (Swin-IR-Diff). The next two baselines are video super-resolution models grounded in vision transformer architecture (VRT) [29] and its recurrent variant (RVRT) [31]. The latter incorporates guided deformable attention for clip alignment, enhancing its temporal modeling capabilities. Lastly, we compare against another video-super-resolution baseline (PSRT) [50] which also relies on the transformer architecture but uses multi-frame attention groups.
@@
We additionally perform ablation studies in three distinct configurations. In the first two ablations, we experiment with the input sequence length. While our proposed model uses a context length of 5 frames, we also conduct experiments with 3 frames and 1 frame (STVD-3 and STVD-1). Note that using a single context frame ablates for the temporal attention block as well. The third ablation (STVD-Single) involves removing the additional input channels (i.e. only providing the model with the low-resolution precipitation sequence) to assess their impact on performance metrics, providing valuable insights into the influence of these supplementary inputs on the model’s overall performance. In summary, our experiments demonstrate that our method outperforms all considered baselines across all metrics considered. Additionally, our ablation studies demonstrate the importance of temporal context and additional climate inputs.
3.1 Dataset
@@
Our data derives from an 11-member initial condition ensemble of 13-month simulations using a global atmosphere model, FV3GFS, run at 25 km resolution, forced by climatological sea surface temperatures and sea ice. The first month of each simulation is discarded to allow the simulations to spin up and meteorologically diverge, effectively providing 11 years of reference data (of which first 10 years are used for training and the last year for validation). FV3GFS, developed by the National Oceanic and Atmospheric Administration (NOAA), is a version of NOAA’s operational global weather forecast model ([67, 7]).
@@
Three-hourly average data were saved from this entire simulation, which was a 25 km horizontal “fine grid” (high-resolution frames). We further 8-fold coarsened the selected fields to a 200 km horizontal “coarse grid" to create paired data ($x_{t},y_{t}$), where $x_{t}$ is the coarse-grid global state, and $y_{t}$ is the corresponding fine-grid global state. Our goal is to apply video downscaling to the coarse-grid precipitation field to obtain temporally smooth fine-grid precipitation estimates that are statistically similar to the true data. This approach is attractive because many fine-grid precipitation features, such as cold fronts and tropical cyclones, are poorly resolved on the coarse grid but are temporally coherent across periods much longer than 3 hours. We use 12 coarse-grid input fields, including precipitation, topography, and horizontal vector wind at various levels. See Appendix A.2 for the list of included atmospheric variables. FV3GFS uses a cubed-sphere grid, where the surface of the globe is divided into six tiles, each of which is covered by an $N\times N$ array of points. Our data fields reflect this structure with $N=48$ for the 200 km coarse grid and $N=384$ for the 25 km fine grid.
@@
The application presented here is a pilot for broader applications of our methodology. Fine-grid simulations are much more computationally expensive than coarse-grid simulations (an 8-fold reduction in grid spacing requires almost 1000x more computation), so a coarse-grid simulation with super-resolved details in desired regions could be highly cost-effective for many applications.
@@
During training, our VSR model randomly selects data at each time from one of the six tiles. This strategy ensures that the model learns from the diverse spatial contexts and weather regimes that produce precipitation in different parts of the world. Post-training, for localized analysis, we selectively sample super-resolved precipitation channels from regions with complex terrain, such as California. These regions can systematically pattern the precipitation on fine scales. This analysis helps us to see how well the super-resolution can learn the time-mean spatial patterns (e.g. precipitation enhancement on the windward side of mountain ranges and lee rain shadows) in the fine-grid reference data.
3.2 Training and Testing Details
@@
Figure 4: Precipitation over two different regions (left: Himalayas; right: Northern California coast, same region as Fig. 2), averaged across a year, for our STVD model and the ground-truth. For each half, the topography of the region is shown in the corresponding top-left whereas the predicted annual average is shown in the corresponding bottom-right. Annually-averaged precipitation is an important indicator of water availability in a region. STVD successfully captures many details of the precipitation that are tied to local topography and are too fine to be resolved the coarse-grid data.
We address the task of downscaling a sequence of precipitation frames with a scale factor of 8 applied to FV3GFS output. Our approach leverages a Spatiotemporal Video Diffusion (STVD) framework, which involves training on 5 consecutive frames, all of which are downscaled jointly.
@@
We optimize our model using Adam [23] with an initial learning rate of $1\times 10^{-4}$ and cosine annealing with decay until $5\times 10^{-7}$ during training, executed on an NVidia RTX A6000 GPU. The diffusion model is trained using v-parametrization [49], with a fixed diffusion depth (N = 1400). Random tiles extracted from the cube-sphere representation of Earth, with dimensions 384 in high-resolution and 48 in low-resolution, are utilized during training. We train for one million steps, requiring approximately 7 days on a single node. We use a batch size of one, apply a logarithmic transformation to precipitation states, and normalize to the range $[-1,1]$.
During testing, we employ DDIM sampling with 30 steps on an Exponential Moving Average (EMA) variant of our model, with a decay rate set to 0.995. Testing is conducted on the full frame size. A single diffusion loss optimises the entire pipeline end-to-end.
3.3 Baseline Models
@@
We compare against several recent and high performing transformer-based video super-resolution models. These models are trained in a deterministic manner, thus allowing us to compare against our generative setup. The first, Video Restoration Transformer (VRT) [29], allows for parallel frame prediction and long-range temporal dependency modeling. The second, recurrent VRT (RVRT) [31], incorporates guided deformable attention for effective clip alignment, enhancing its temporal modeling capabilities. PSRT [50] removes the alignment module and modifies the attention window.
@@
In addition, to assess the benefits of multi-frame downscaling, we compare to Swin-IR [30]. This model is a popular image super-resolution model that harnesses Swin Transformer blocks. Swin-IR is, however, still trained in a supervised fashion. Thus, as a generative baseline, we compare to Swin-IR-Diff. This model generates a deterministic prediction using Swin-IR [30], followed by modeling of a stochastic residual via diffusion. This baseline is inspired by concurrent work on single-image radar-reflectivity downscaling [35], where a UNet is used instead of Swin-IR. See Appendix A.2 for details.
3.4 Evaluation Metrics
@@
We evaluate our model differently than for standard vision tasks. In addition to the Mean Square Error (MSE), which measures the average squared difference between predicted and actual values but does not capture the full distributional information, a more meaningful comparison includes several distribution-level metrics. One such important metric is the Continuous Ranked Probability Score (CRPS) [5, 56], used to assess the discrepancy between the predicted cumulative distribution function and the observed data. We compute CRPS over 10 stochastic realisations of our prediction for a give low-resolution sequence.
@@
Furthermore, given the distinctive distribution of the precipitation climate state, characterized by a light-tailed exponential nature, it is important to ensure that the application of downscaling does not induce significant alterations in the distribution of precipitation rates. This gives rise to two additional metrics. First, we compute the Earth Mover Distance [44] to quantify the agreement between the target and predicted global precipitation distribution, which is strongly affected by high-resolution details. Second, focusing on tail events and extreme precipitation, we consider the 99.999th percentile error (PE) in the global precipitation distribution, providing a nuanced understanding of the model’s performance on rare and extreme precipitation events.
Table 1: Quantitative comparison between our STVD method and the various competitive baselines. EMD represents the Earth-Mover Distance, and PE denotes the 99.999th percentile error. Overall, our proposed method (STVD) outperforms the baselines across all metrics. The exclusion of additional side information (STVD-single) or decrement in context length (STVD-3 and STVD-1) appreciably degrades performance.
CRPS MSE EMD PE
$(10^{-5})$ $(10^{-8})$ $(10^{-6})$ $(10^{-3})$
STVD (ours) $\mathbf{1.85}$ $\mathbf{0.59}$ $\mathbf{2.49}$ $\mathbf{1.2}$
PSRT [50] $2.15$ $0.66$ $4.21$ $3.8$
RVRT [31] $3.55$ $1.73$ $4.33$ $3.6$
VRT [29] $3.58$ $1.74$ $4.61$ $4.0$
Swin-IR-Diff [35, 46] $2.29$ $1.94$ $6.38$ $4.4$
Swin-IR [30] $2.36$ $2.29$ $17.40$ $23.40$
[…]
STVD-single (ablation) $1.81$ $0.62$ $4.64$ $2.3$
STVD-3 (ablation) $1.96$ $0.68$ $4.94$ $2.6$
STVD-1 (ablation) $2.05$ $0.72$ $7.19$ $4.1$
3.5 Qualitative and Quantitative Analysis
@@
Tab. 1 presents a comprehensive quantitative evaluation comparing our method with state-of-the-art baselines and ablations. Our proposed model (STVD) exhibits strong performance across all metrics, outperforming all considered baselines. We highlight the distributional characteristics in Fig. 3. Swin-IR overestimates precipitation, whereas all other baselines underestimate precipitation. This is a highly undesirable property of a downscaling model, as poor performance on rare and extreme precipitation events can adversely affect marginal disaster mitigation policies. In contrast, our method closely matches the precipitation distribution, as measured by the 99.999th percentile error and EMD.
@@
The use of only precipitation as an input (STVD-single) results in slightly worse performance across all metrics, indicating that the additional inputs are indeed predictive. In contrast, our ablation model STVD-1, lacking the information from the full sequence, shows a significant difference, emphasizing the importance of temporal attention in our approach.
@@
Fig. 2 depicts the performance of our model compared to other baselines on example of a precipitation feature interacting with mountainous terrain, using a three-hourly succession of plots. A cold front (linear feature in the left-most column of plots) impinges on northern California, producing enhanced precipitation first over the coastal foothills and then over the northwest-southeast ridge of the Sierra mountain range in the lower right corner of the images. Our model generates high quality results which preserves most patterns with a high degree of similarity. RVRT produces slightly more diffuse precipitation features, while Swin-IR produces slightly more pixelated precipitation features.
@@
Fig. 4 shows annually-averaged precipitation from the patch in Fig. 2. Accurately capturing the fine-grid structure of time-mean precipitation is an informative metric and important for assessing long-term water availability. Our method (which includes fine-grid topography as a training input) is remarkably successful in replicating the ground truth, including the strength and narrow spatial structure of the bands of high precipitation along the Northern California coastal mountains and the Sierras. These features are not resolved by the coarse-grid inputs to the super-resolution. Appendix A.3 has additional samples.

### v3

@@ 2.1.3 Loss Function
To train our model, we use the angular parametrization suggested by [54]. Specifically, this results in the diffusion loss of the form
@@
where $\epsilon\sim\mathcal{N}(0,I)$, $n$ is sampled uniformly from $\{1,2,\dots,N\}$, and the sequences $\mathbf{x}^{0:T}$, $\mathbf{y}^{0:T}$ are sampled from the training distribution. Here, $\mathbf{c}=(\mathbf{x}^{0:T},\overline{\mathbf{y}}^{0:T})$ where $\bar{\mathbf{y}}^{0:T}=\mu_{\phi}\left(\mathbf{x}^{0:T}\right)$. The scalars $\alpha_{n}^{2}=\prod_{i=1}^{n}(1-\beta_{i})$ and $\sigma_{n}^{2}=1-\alpha_{n}^{2}$ are used to define $\mathbf{v}\equiv\alpha_{n}\epsilon-\sigma_{n}\mathbf{r}_{0}^{0:T}$. Training and inference are concurrent across multiple frames due to spatio-temporal attention. Alg. 1 and 2 demonstrate the training and sampling strategy under the angular parametrization. We use DDIM sampling [57] to generate frame residuals with fewer diffusion steps.
@@ 2.1.4 Network Architecture
Both the downscaler and the conditional diffusion model employ a UNet backbone with similar architectures and key adaptations to the attention mechanism (see A.1). The downscaler takes the multi-channel input frames ($\mathbf{x}^{0:T}$), yielding an initial estimate ($\bar{\mathbf{y}}^{0:T}$). The diffusion UNet conditions on diffusion step $n$ and concatenates feature maps from the downscaler with its own. The concatenated input to the diffusion UNet (${\mathbf{x}}^{0:T}$, $\bar{\mathbf{y}}^{0:T}$, and $\mathbf{r}^{0:T}_{n}$), along with the conditioning variables (diffusion step $n$ and the feature maps from downscaler), yields the output $\mathbf{v}$.
Computing full attention for temporal coherence across the entire video data cube is very expensive for processing long sequences or high-resolution inputs. To optimize efficiency, we decouple attention between spatial and temporal dimensions, use a linear variant of self-attention [24] for non-bottleneck layers (where the effective number of “tokens” for attention is relatively large), focus spatial attention on localized patches (instead of the entire feature map, which could be wasteful), and calculate per-channel temporal attention in large spatial dimensions (namely, the ultimate and penultimate expansion and contraction layers of UNet). These modifications dramatically reduce the time complexity and memory footprint of these transformer blocks.
@@ 3 Experiments
We conduct a comprehensive evaluation of our proposed method, SpatioTemporal Video Diffusion (STVD), against six contemporary state-of-the-art baselines. The first two baselines are image super-resolution models based on the Swin Vision Transformer (Swin-IR)[31] and its residual diffusion variant (Swin-IR-Diff). The next two baselines are video super-resolution models grounded in vision transformer architecture (VRT) [30] and its recurrent variant (RVRT) [32]. The latter incorporates guided deformable attention for clip alignment, enhancing its temporal modeling capabilities. We compare against another video-super-resolution baseline (PSRT) [55] which also relies on the transformer architecture but uses multi-frame attention groups. Finally, we compare against a video diffusion baseline (VDM) [19].
@@
We perform ablation studies in three configurations. In the first two, we experiment with the input sequence length. While our proposed model uses a context length of 5 frames, we also conduct experiments with 3 frames and 1 frame (STVD-3 and STVD-1). Note that using a single context frame ablates for the temporal attention block as well. The third ablation (STVD-Single) involves removing the additional input channels (i.e. only providing the model with the low-resolution precipitation sequence) to assess their impact on performance metrics. In summary, our experiments demonstrate that our method outperforms all baselines across all metrics considered. Additionally, our ablation studies highlight the importance of temporal context and additional climate inputs. [Uncaptioned image] Figure 4: Distributions of the fine-grid three-hourly average precipitation, for all gridpoints around the globe. The Swin-IR baseline overestimates large precipitation events, whereas all other baselines underestimate key extreme and rare precipitation events. Our model aligns best with the fine-grid ground truth than any the other model. This is also evident with the the EMD and PE metrics discussed in Tab. 1 and Sec. 3. Table 1: Quantitative comparison between our method and other competitive baselines. EMD represents the Earth-Mover Distance, PE denotes the 99.999th percentile error and SAE is the spatial-autocorrelation error. Overall, our proposed method (STVD) outperforms the baselines across all metrics. The exclusion of additional side information (STVD-single) or decrement in context length (STVD-3 and STVD-1) appreciably degrades performance. CRPS MSE EMD PE SAE $(10^{-5})$ $(10^{-8})$ $(10^{-6})$ $(10^{-3})$ $(10^{-6})$ STVD (ours) $\mathbf{1.85}$ $\mathbf{0.59}$ $\mathbf{2.49}$ $\mathbf{1.2}$ $\mathbf{4.00}$ PSRT [55] $2.15$ $0.66$ $4.21$ $3.8$ $6.24$ RVRT [32] $3.55$ $1.73$ $4.33$ $3.6$ $7.39$ VRT [30] $3.58$ $1.74$ $4.61$ $4.0$ $7.39$ Swin-IR-Diff [37, 51] $2.29$ $1.94$ $6.38$ $4.4$ $7.70$ VDM [19] $2.21$ $0.73$ $12.70$ $6.4$ $8.84$ Swin-IR [31] $2.36$ $2.29$ $17.40$ $23.40$ $18.9$hdashline STVD-single (ablation) $1.81$ $0.62$ $4.64$ $2.3$ $6.09$ STVD-3 (ablation) $1.96$ $0.68$ $4.94$ $2.6$ $4.99$ STVD-1 (ablation) $2.05$ $0.72$ $7.19$ $4.1$ $6.87$
Dataset
@@
Our dataset derives from an 11-member initial condition ensemble of 13-month simulations using a global atmosphere model, FV3GFS, run at 25 km resolution and forced by climatological sea surface temperatures and sea ice. The first month of each simulation is discarded to allow the simulations to spin up and meteorologically diverge, effectively providing 11 years of reference data (of which first 10 years are used for training and the last year for validation). FV3GFS, developed by the National Oceanic and Atmospheric Administration (NOAA), is a version of NOAA’s operational global weather forecast model ([72, 9]).
@@
Three-hourly average data were saved from this entire simulation, which used a 25 km horizontal “fine grid”. We further coarsened the selected fields by a factor of 8 to create a 200 km horizontal “coarse grid", resulting in paired data ($x_{t},y_{t}$), where $x_{t}$ is the coarse-grid global state and $y_{t}$ is the corresponding fine-grid global state. Our goal is to apply video downscaling to the coarse-grid precipitation field to obtain temporally smooth fine-grid precipitation estimates that are statistically similar to the true data. This approach is attractive because many fine-grid precipitation features, such as cold fronts and tropical cyclones, are poorly resolved on the coarse grid but are temporally coherent across periods much longer than 3 hours. We use 12 coarse-grid input fields, including precipitation, topography, and horizontal vector wind at various levels. See A.2 for the list of included atmospheric variables. FV3GFS uses a cubed-sphere grid, where the surface of the globe is divided into six tiles, each of which is covered by an $S\times S$ array of points. Our data fields reflect this structure with $S=48$ for the 200 km coarse grid and $S=384$ for the 25 km fine grid.
@@
The application presented here serves as a pilot for broader uses of our methodology. Fine-grid simulations are significantly more computationally expensive than coarse-grid simulations (an 8-fold reduction in grid spacing requires almost 1000x more computation), so a coarse-grid simulation with super-resolved details in desired regions could be highly cost-effective for many applications.
@@
During training, our model randomly selects data from one of the six tiles. This strategy ensures that the model learns from the diverse spatial contexts and weather regimes that produce precipitation worldwide. Post-training, for localized analysis, we selectively sample super-resolved precipitation channels from regions with complex terrain, such as California (Fig. 5). These regions can systematically pattern the precipitation on fine scales. This analysis helps us to see how well the super-resolution can learn the time-mean spatial patterns (e.g. precipitation enhancement on the windward side of mountain ranges and lee rain shadows) in the fine-grid reference data.
Training and Testing Details
@@
Figure 5: Precipitation over two regions (left: Himalayas; right: Northern California coast, same region as Fig. 2), averaged across a year, for our STVD model and the ground-truth. For each half, the topography of the region is shown in the corresponding top-left whereas the predicted annual average is shown in the corresponding bottom-right. Annually-averaged precipitation is an important indicator of water availability in a region. STVD successfully captures many details of the precipitation that are tied to local topography and are too fine to be resolved the coarse-grid data.
@@
We downscale a sequence of precipitation frames from FV3GFS output by a factor of 8. Our approach (STVD) trains on 5 consecutive frames that are downscaled jointly. We optimize our model end-to-end with a single diffusion loss using Adam [26] with an initial learning rate of $1\times 10^{-4}$, decaying to $5\times 10^{-7}$ with cosine annealing during training, executed on an NVidia RTX A6000 GPU. The diffusion model is trained using v-parametrization [54], with a fixed diffusion depth (N = 1400). Random tiles extracted from the cube-sphere representation of Earth, with dimensions 384 in high-resolution and 48 in low-resolution, are used during training. We train for one million steps, requiring approximately 7 days on a single node (slightly lesser for ablations). We use a batch size of one, apply a logarithmic transformation to precipitation states, and normalize to the range $[-1,1]$. During testing, we employ DDIM sampling with 30 steps on an Exponential Moving Average (EMA) variant of our model (for full frame size), with a decay rate of 0.995.
Baseline Models
@@
We compare our generative setup against several recent high-performing transformer-based video super-resolution models. These models are trained deterministically. The first, Video Restoration Transformer (VRT) [30], allows for parallel frame prediction and long-range temporal dependency modeling. The second, recurrent VRT (RVRT) [32], incorporates guided deformable attention for effective clip alignment, enhancing its temporal modeling capabilities. The third, PSRT [55], removes the alignment module and modifies the attention window. We also compare against the recent Video Diffusion Model (VDM) [19] which employs global quadratic attention.
@@
To assess the benefits of multi-frame downscaling, we compare with Swin-IR [31], a popular image super-resolution model that harnesses Swin Transformer blocks. However, Swin-IR is trained in a supervised fashion. Thus, as a generative baseline, we compare to Swin-IR-Diff. This model generates a deterministic prediction using Swin-IR [31], followed by modeling a stochastic residual using diffusion. This baseline is inspired by concurrent work on single-image radar-reflectivity downscaling [37], where a UNet is used instead of Swin-IR. See A.2 for details.
Evaluation Metrics
@@
We evaluate our model differently from standard vision tasks. In addition to the Mean Square Error (MSE), which measures the average squared difference between predicted and actual values but lacks full distributional information, we use several distribution-level metrics for a more meaningful comparison. One such metric is the Continuous Ranked Probability Score (CRPS) [7, 61], which assesses the discrepancy between the predicted cumulative distribution function and the observed data. We compute CRPS over 10 stochastic realizations of our predictions.
@@
Furthermore, given the distinctive light-tailed exponential distribution of the precipitation climate state, it is crucial to ensure that downscaling does not significantly alter the distribution of precipitation rates. This necessitates two additional metrics. First, we compute the Earth Mover Distance [49] to quantify the agreement between the target and predicted global precipitation distributions, which are strongly affected by high-resolution details. Second, we focus on tail events and extreme precipitation by considering the 99.999th percentile error (PE), providing a nuanced understanding of the model’s performance on rare and extreme precipitation events.
To further assess the spatial fidelity of our downscaling approach, we use the Spatial Autocorrelation Error (SAE) [63]. This metric calculates the mean absolute error between the spatial autocorrelation of the predictions and ground truth. Low SAE ensures that the spatial patterns and the fine structure in precipitation data are preserved during downscaling, which is critical for accurate climate modeling.
Qualitative and Quantitative Analysis
@@
Tab. 1 provides a quantitative evaluation comparing our method with state-of-the-art baselines and ablations. Our model (STVD) performs strongly across all metrics, outperforming all baselines. We highlight the distributional characteristics in Fig. 4. Swin-IR overestimates precipitation, while all other baselines underestimate it. This discrepancy is undesirable, as poor performance on rare and extreme precipitation events can negatively imapct disaster mitigation policies. In contrast, our method closely matches the precipitation distribution, as measured by PE and EMD.
@@
Using only precipitation as an input (STVD-single) results in slightly worse performance across all metrics, indicating the predictive value of additional inputs. In contrast, our ablation model STVD-1, which lacks full sequence information, performs significantly worse, highlighting the importance of temporal attention in our approach.
@@
Figs. 2, 9 and 10 depict the performance of our model compared to other baselines on examples of a precipitation feature interacting with mountainous terrain. Our model generates high quality results which preserves most patterns with a high degree of similarity. PSRT and RVRT produce slightly more diffuse precipitation features, while Swin-IR produces slightly more pixelated features.
@@
Fig. 5 shows annually-averaged precipitation from the patches in Figs. 2 and 10. Accurately capturing the fine-grid structure of time-mean precipitation is crucial for assessing long-term water availability. Our method (which includes fine-grid topography as a training input) effectively replicates the ground truth. This includes the strength and narrow spatial structure of high precipitation bands along the Northern California coastal mountains and the Sierras. These features are not resolved by the coarse-grid inputs to the super-resolution. See A.3 for full-page high-resolution samples and spectral analysis. Fig. 11 reveals that the spectra for baselines decay more rapidly than for STVD.
Realism-Distortion Tradeoff
Distortion metrics such as MSE often conflict with perceptual quality, wherein reducing distortion typically degrades perceptual realism [6]. In our context, this tradeoff translates to balancing MSE and PE. While MSE captures the average accuracy of predictions, PE represents the model’s ability to reproduce extreme events, thereby serving as a proxy for realism. Realism in climate modeling refers to the accurate representation of extreme weather patterns, which are crucial for applications like flood forecasting and disaster mitigation. PE is a distributional criterion that effectively captures these tail events, offering a robust measure of realism. Fig. 3 illustrates this tradeoff, with darker colors corresponding to fewer STVD sampling steps. As the number of sampling steps increases, MSE tends to rise slightly, but PE decreases significantly. Depending on the application, this tradeoff may potentially be exploited by practitioners.